Five jobs are ready for processing at time zero through a workstation. The details of the processing time and the due date of completion is provided in the following table.

Job

Processing Time (days)

Due Date (days)

A

5

11

B

6

16

C

3

19

D

4

10

E

2

9



Using the earliest due date (EDD) rule, the average lateness for the five jobs is:

a. more than -6 but less than or equal to -2 days.
b. more than -2 but less than or equal to 2 days.
c. more than 2 but less than or equal to 6 days.
d. more than 4 but less than or equal to 8 days.
Answer #1

The Jobs will be scheduled as earliest due date to latest due date

EDD Rule

Jobs

Processing Time (days)

Completion of job (Cumulative)

Due Date (days)

Tardiness or lateness (if completion time is more than due date)

E

2

2

9

0

D

4

6

10

0

A

5

11

11

0

B

6

17

16

1

C

3

20

19

1

The average lateness for the five jobs = (Total late days/No. of Jobs)

= (1 + 1)/ 5 = 2/5 = 0.4 days

Therefore correct answer is option b. more than -2 but less than or equal to 2 days.